OpenStudy (anonymous):

i have a question.... Input the equation of the given line in standard form. The line including (3, 1) and (-2, 3).

hartnn (hartnn):

can you find the slope since 2 points are given ?

OpenStudy (anonymous):

yes, 3-1=2 -2-3=-5 slope = 2/-5

hartnn (hartnn):

correct! now use the slope point form y-y1 = m (x-x1) y1 =1 x1=3 m=-2/5 just plug in values!

OpenStudy (anonymous):

y-1=-2/5(x-3)

hartnn (hartnn):

yes, now multiply 5 on both sides to eliminate fraction..

OpenStudy (anonymous):

5/y-1=-2(x-3) like this>>>

hartnn (hartnn):

\( \Large 5 \times ( y-1) = (-2/\cancel 5 \times \cancel 5)(x-3) \\ \Large 5(y-1)=-2 (x-3) \) like this...

OpenStudy (anonymous):

5y-5=-2x+6

hartnn (hartnn):

yep, add 2x on both sides

OpenStudy (anonymous):

2x+5y-5=6

hartnn (hartnn):

finally add 5

OpenStudy (anonymous):

5x+5y=11

hartnn (hartnn):

2x+5y =11 you mean ?

OpenStudy (anonymous):

oh right ....thankx 4 the help...

hartnn (hartnn):

welcome ^_^